Ester= Isoamyl acetate (banana)

Molar mass of banana= 130.19

density= 0.876 g/ml

5.26 g ester product.

Alcohol= isoamyl alcohol

density= 0.81 g/ml

Molar mass= 88.148 g/mol

Quantity= 5.4 ml

Carboxylic acid= acetic acid

density= 1.05 g/ml

Molar mass= 60.052 g/mol

Quantity= 11.4 ml

and 1 ml of concentrated H2SO4 (1.8 g, 0.018 mole) was also used in the procedure.

Calculate theoretical yield and % yield .
Answer #1

Mass of Alconal= 5-4m) X0.819|m) = 4.3748 . moles of Alcohol = 4.374 188-1488/ml = 0.0496 me mast of carboxylic Acid = ilumi